# Holographic entanglement negativity and thermodynamics in backreacted AdS black hole

### v1

@@ 3.2 Entanglement wedge cross section and reflected entropy
Unlike HEE, the Entanglement Wedge Cross Section (EWCS) required two non-overlapping intervals (say $A_{1}$ and $A_{2}$). It is a geometric quantity associated with the minimal bulk surface that divides the entanglement wedge of two boundary sub regions $A_{1}$ and $A_{2}$ into two parts homologous to each of them. The EWCS has emerged as an important mixed state correlation measure in holography, conjectured to be dual to the entanglement of purification [Takayanagi:2017knl, Nguyen:2017yqw], odd entanglement entropy [Tamaoka:2018ned], entanglement negativity [Kudler-Flam:2018qjo, Kusuki:2019zsp, KumarBasak:2020eia, KumarBasak:2021lwm] and widely studied at various occasions, see [BabaeiVelni:2019pkw, Jokela:2019ebz, Jeong:2019xdr, Jain:2020rbb, Basu:2022nds, Jain:2022hxl, Jain:2022csf, Jiang:2024ijx]. In [Dutta:2019gen] authors proposed a relation between reflected entropy and EWCS via the following relation
@@
The EWCS between $A_{1}$ and $A_{2}$ is defined as the minimal area surface $\Gamma_{A_{1}A_{2}}$ that connects the two RT surfaces $\gamma_{A_{1}}$ and $\gamma_{A_{2}}$ bounding the entanglement wedge of $A_{1}\cup A_{2}$. Its area is given by
$E_{W}(A_{1},A_{2})=\frac{\mathcal{A}(\Gamma_{A_{1}A_{2}})}{4G_{N}^{d+1}}=\frac{R^{d-1}L^{d-2}}{4G_{N}}\int_{z_{D}}^{z_{t}}\frac{dz}{z^{d-1}\sqrt{h(z)}},$ (3.17)
@@
$\begin{split}E_{W}^{low}&=\frac{R^{d-1}L^{d-2}}{4G_{N}}\Bigg[K_{0}\!\left(\frac{1}{D^{d-2}}-\frac{1}{(D+2l)^{d-2}}\right)+K_{1}\,\rho\,l\!\left(\frac{4\pi T_{f}}{d}\right)^{\!d-1}\\ &\qquad\qquad\quad-\,K_{2}(1-\rho)\,l(l+D)\!\left(\frac{4\pi T_{f}}{d}\right)^{\!d}+\mathcal{O}\big((T_{f}l)^{2d-2}\big)\Bigg],\end{split}$ (3.18)
where the coefficients $K_{0},K_{1},K_{2}$ depend only on the dimension $d$ and are given explicitly in [Chakrabortty:2020ptb]. The first term corresponds to the vacuum contribution, while the $\rho$-dependent terms encode the leading and subleading backreaction effects. Notably, the correction proportional to $\rho T_{f}^{d-1}$ enhances the EWCS, indicating stronger mixed-state correlations in the presence of quark backreaction at low temperature.
@@
with $\mathcal{C}$ is a function of $\rho$ and one can read form [Chakrabortty:2020ptb]. In this regime, EWCS scales as $T_{f}^{d-2}$, and the backreaction enhances its magnitude monotonically. In the limit $\rho\to 0$, both (3.18) and (3.19) reduce to the AdS–Schwarzschild results, providing a nontrivial consistency check.
@@
In holographic theories, the reflected entropy $S_{R}(A_{1},A_{2})$ is proposed to be related to the EWCS via (3.15) [Dutta:2019gen]. This relation is supported by explicit checks in static and time-dependent AdS3/CFT2 setups as well as tensor-network constructions [Nguyen:2017yqw, Takayanagi:2017knl, Umemoto:2018jpc]. Consequently, the low and high temperature expressions for reflected entropy follow directly from (3.18) and (3.19) by an overall factor of two.
@@
Although reflected entropy is geometrically determined by EWCS at the classical level, its interpretation as the entanglement entropy of a canonical purification provides additional insight beyond the purely geometric picture. In particular, while EWCS diagnoses the connectivity of the entanglement wedge, $S_{R}$ quantifies the total mixed-state entanglement shared between the subsystems, including correlations inaccessible to entanglement negativity. In the backreacted AdS black hole background, this distinction becomes important: EWCS and consequently $S_{R}$ is enhanced by backreaction at high temperature, whereas HEN exhibits a qualitatively different behavior in the low-temperature regime. This contrast highlights the complementary roles of EWCS, reflected entropy, and negativity in characterizing the entanglement structure of strongly coupled plasmas.
@@ 4 Entanglement thermodynamics
$T_{H}=\frac{1}{4\pi z_{H}}\left(3-\frac{bz_{H}^{2}}{R^{2}}\right).$ (4.4)
@@
${\cal A}$ $=\frac{2LR^{2}}{z_{t}}\int_{0}^{1}\frac{}{u^{2}\sqrt{1-u^{4}}}\left(1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}+\frac{bz_{H}^{2}}{R^{2}}\left(\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}-\frac{u^{2}z_{t}^{2}}{z_{H}^{2}}\right)\right)^{-\frac{1}{2}}du.$ (4.7)
We utilize the subsystem length and area formulas given in eq.˜4.6 and eq.˜4.7 to investigate the thermal and backreaction effects on HEE. For a field theory dual to a deformed AdS4 spacetime, the space of states is is characterized by both $T$ (via $z_{H}$) and $b$. This is in contrast to the Schwarzschild black hole in the $AdS_{4}$ bulk scenario where $T$ is the unique scale controlling the high and low temperature limits of the entropy. In contrast, for the deformed black hole, the backreaction parameter $b$ also plays a role, as is evident from the requirement of temperature positivity in the bulk which follows from eq.˜4.4 as
@@ 4.1 Zero temperature
Thus, for deformed black holes zero temperature, the small backreaction regime also corresponds to a large horizon radius $(l^{2}b\leq\frac{3R^{2}l^{2}}{z_{H}^{2}}<<1)$. With this condition, the lapse function from eq.˜4.2 simplifies
@@
${\cal A}$ $=\frac{2LR^{2}}{z_{t}}\int_{0}^{1}\frac{}{u^{2}\sqrt{1-u^{4}}}\left(1-\frac{3u^{2}z_{t}^{2}}{z_{H}^{2}}+\frac{2u^{3}z_{t}^{3}}{z_{H}^{3}}\right)^{-\frac{1}{2}}du.$ (4.12)
@@
The first term in the expression for $\mathcal{A}$ corresponds to the divergent contribution which is same as of pure AdS4. We regularize the integral by imposing a UV cutoff $\epsilon=z_{b}/z_{t}$ and incorporate a counterterm $(-2LR^{2}_{AdS})/z_{b}$ to subtract the divergence. This procedure gives the following finite value for the area as
@@
$S_{A}^{finite}=\frac{1}{4G_{N}^{4}}\frac{LR^{2}}{l}\left[\frac{\pi\Gamma\left(\frac{3}{4}\right)\Gamma\left(-\frac{1}{4}\right)}{\Gamma\left(\frac{1}{4}\right)^{2}}+\frac{\pi\sqrt{2}\Gamma\left(\frac{5}{4}\right)b\,l^{2}}{3R^{2}_{AdS}\Gamma\left(\frac{3}{4}\right)^{3}}+\frac{16\pi\sqrt{2}l^{3}\Gamma\left(\frac{5}{4}\right)}{\Gamma\left(-\frac{1}{4}\right)^{3}z_{H}^{3}}+\mathcal{O}(\frac{l}{z_{H}})^{4}\right].$ (4.18)
Therefore, in the zero-temperature case, the HEE receives corrections that are linear and quadratic in $l$ and $l^{2}$. Notably, the emergence of a term linear in $l$ distinguishes this geometry from the extremal RN black hole case [Chaturvedi:2016kbk] where it is absent. By substituting the zero temperature mass $m_{0}=-R^{4}/z_{H}^{3}$ obtained from eq.˜4.3 and eq.˜4.9 into above eq.˜4.18, the renormalized HEE can be expressed in the following compact form
@@ 4.2 Low temperature regime
We now turn to the analysis of the subsystem $A$ in the boundary field theory dual to a deformed $AdS_{4}$ black hole with small backreaction at low temperature. As follows from eq.˜4.8, when both the temperature and the backreaction are small, the horizon radius $z_{H}$ becomes large. In this limit, one has $bz_{H}^{2}/R^{2}\sim 1$ and $z_{H}$ sufficiently large such that $z_{t}\ll z_{H}$. Under these conditions, the lapse function $h(u)$ takes the form
@@ 4.3 First law of entanglement thermodynamics
$\Delta E_{A}=\int_{A}dx^{2}\ T_{tt}^{T_{H}\neq 0}-\int_{A}dx^{2}\ T_{tt}^{T_{H}=0}=\frac{Ll}{4\pi G_{N}^{4}R^{2}}(m-m_{0}),$ (4.32)
where $T_{tt}$ denotes the time component of the boundary stress-energy tensor and can be obtained using the holographic prescription of [Balasubramanian:1999re]. Now substituting the results from eq.˜4.31 and eq.˜4.32 into the first law like relation eq.˜4.30, the entanglement temperature is found to be
@@
We observe that the entanglement temperature $T_{\mathrm{ent}}$ is proportional to the inverse of the subsystem size. This universal scaling behavior is consistent with the entanglement thermodynamics of the asymptotically AdS geometries as discussed in [Bhattacharya:2012mi]. This suggests that the small backreaction does not alter the fundamental scaling of the first law and the HEE continues to satisfy a well defined thermodynamic relation in the perturbative regime.
@@ 4.4 High temperature regime
In this section, we analyze the high temperature behavior of the HEE for a subsystem $A$ when the bulk deformed black hole has a small backreaction and is in the high-temperature regime. From eq.˜4.8, it follows that in the limit of high temperature ($l/z_{H}>>1$) and small backreaction, the parameter $bz_{H}^{2}/3R^{2}$ becomes much smaller than unity i.e $\frac{bz_{H}^{2}}{3R^{2}}\ll 1$. By introducing the small parameter $\delta=bz_{H}^{2}/3R^{2}$, we can Taylor expand the lapse function $h(u)^{-1/2}$ around $\delta=0$ yielding
@@
$\mathcal{A}$ $=\frac{2LR^{2}}{z_{t}}\int_{0}^{1}\frac{1}{u^{2}\sqrt{1-u^{4}}}\left(\frac{1}{\sqrt{1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}}}+\frac{3\delta u^{2}z_{t}^{2}\left(1-\frac{uz_{t}}{z_{H}}\right)}{2z_{H}^{2}\left(1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}\right){}^{3/2}}\right)\,du$ (4.36)
@@
As described in[Hubeny:2012ry], for a static asymptotically AdS black hole, the extremal surface cannot penetrate the event horizon, necessitating that $z_{t}<z_{H}$. In the high temperature limit, however, the turning point approaches the horizon such that $z_{t}\to z_{H}$. By introducing a small positive parameter $\epsilon\ll 1$, we can parameterize this proximity as $z_{t}=(1-\epsilon)z_{H}$. Taking the limit $\epsilon\to 0$ in eq.˜4.39, the series converges to the following compact leading order expression as follows
@@
In evaluating the area integral given in eq.˜4.36, divergent part is confined to the first term. Following [Fischler:2012ca], we isolate the divergence and express the remaining finite contribution as
${\cal A}^{finite}=\frac{2LR^{2}}{z_{t}}\int_{\frac{z_{b}}{r_{t}}}^{1}\frac{1}{u^{2}\sqrt{1-u^{4}}}\left(\frac{1}{\sqrt{1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}}}+\frac{3\delta u^{2}z_{t}^{2}\left(1-\frac{uz_{t}}{z_{H}}\right)}{2z_{H}^{2}\left(1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}\right){}^{3/2}}\right)\,du-\frac{2L}{z_{b}}.$ (4.42)
@@
We can now use eq.˜4.38 in above eq.˜4.43 to simplify the extremal area as
@@
Here the term $s=\frac{R^{2}}{4G_{N}^{4}z_{H}^{2}}$ corresponds to the entropy density of the black hole as can be seen from from the eq.˜2.13. In eq.˜4.47, the first term follows a volume law scaling (proportional to the subsystem area in the $(2+1)$ dimensional dual field theory) and grows with temperature indicating that it represents the thermal part of the entropy. The subleading terms scale with the length of the entangling surface separating subsystem $A$ from its complement and thus represents entanglement contributions at high temperatures. Notably, the $\epsilon$-corrections exhibit an exponential decay with temperature analogous to the Schwarzschild-AdS case although they are modified by a small $\delta$ term arising from the backreaction.
@@
Entanglement negativity is a well established quantum information theoretic measure that captures the distillable quantum correlations in mixed states, making it particularly suitable for thermal or finite-density systems where entanglement entropy alone fails to isolate genuine quantum correlations. Defined via the partial transpose of the reduced density matrix, negativity provides a computable upper bound on the distillable entanglement. Within the AdS/CFT framework, a holographic prescription for entanglement negativity in higher dimensions has been proposed, wherein suitable linear combinations of Ryu–Takayanagi surfaces encode the negativity of boundary mixed states [Chaturvedi:2016rft, Jain:2017xsu, Jain:2018bai, KumarBasak:2020viv, Mondal:2021kzj, Afrasiar:2021hld]. This construction allows one to extend the study of mixed state entanglement to strongly coupled holographic plasmas, and in what follows we analyze its behavior in the deformed AdS black hole background to compare with the trends observed for HEN.

### v2

@@ 3.2 Entanglement wedge cross section and reflected entropy
Unlike HEE, the Entanglement Wedge Cross Section (EWCS) required two non-overlapping intervals (say $A_{1}$ and $A_{2}$). It is a geometric quantity associated with the minimal bulk surface that divides the entanglement wedge of two boundary sub regions $A_{1}$ and $A_{2}$ into two parts homologous to each of them. The EWCS has emerged as an important mixed state correlation measure in holography, conjectured to be dual to the entanglement of purification [78, 67], odd entanglement entropy [79], entanglement negativity [50, 56, 52, 53] and widely studied at various occasions, see [3, 44, 42, 36, 5, 34, 35, 43]. In [26] authors proposed a relation between reflected entropy and EWCS via the following relation
@@
The EWCS between $A_{1}$ and $A_{2}$ is defined as the minimal area surface $\Gamma_{A_{1}\cup A_{2}}$ that connects the two RT surfaces $\gamma_{A_{1}}$ and $\gamma_{A_{2}}$ bounding the entanglement wedge of $A_{1}\cup A_{2}$. Its area is given by
$E_{W}(A_{1},A_{2})=\frac{\mathcal{A}(\Gamma_{A_{1}\cup A_{2}})}{4G_{N}^{d+1}}=\frac{R^{d-1}L^{d-2}}{4G_{N}^{d+1}}\int_{z_{D}}^{z_{t}}\frac{dz}{z^{d-1}\sqrt{h(z)}},$ (3.17)
@@
$\begin{split}E_{W}^{low}&=\frac{R^{d-1}L^{d-2}}{4G_{N}^{d+1}}\Bigg[K_{0}\!\left(\frac{1}{D^{d-2}}-\frac{1}{(D+2l)^{d-2}}\right)+K_{1}\,\rho\,l\!\left(\frac{4\pi T_{f}}{d}\right)^{\!d-1}\\ &\qquad\qquad\quad-\,K_{2}(1-\rho)\,l(l+D)\!\left(\frac{4\pi T_{f}}{d}\right)^{\!d}+\mathcal{O}\big((T_{f}l)^{2d-2}\big)\Bigg],\end{split}$ (3.18)
where the coefficients $K_{0},K_{1},K_{2}$ depend only on the dimension $d$ and are given explicitly in [14]. The first term corresponds to the vacuum contribution, while the $\rho$-dependent terms encode the leading and subleading backreaction effects. Notably, the correction proportional to $\rho T_{f}^{d-1}$ enhances the EWCS, indicating stronger mixed-state correlations in the presence of quark backreaction at low temperature.
@@
with $\mathcal{C}$ is a function of $\rho$ and one can read form [14]. In this regime, EWCS scales as $T_{f}^{d-2}$, and the backreaction enhances its magnitude monotonically. In the limit $\rho\to 0$, both (3.18) and (3.19) reduce to the AdS–Schwarzschild results, providing a nontrivial consistency check.
@@
In holographic theories, the reflected entropy $S_{R}(A_{1},A_{2})$ is proposed to be related to the EWCS via (3.15) [26]. This relation is supported by explicit checks in static and time-dependent AdS3/CFT2 setups as well as tensor-network constructions [67, 78, 80]. Consequently, the low and high temperature expressions for reflected entropy follow directly from (3.18) and (3.19) by an overall factor of two.
@@
Although reflected entropy is geometrically determined by EWCS at the classical level, its interpretation as the entanglement entropy of a canonical purification provides additional insight beyond the purely geometric picture. In particular, while EWCS diagnoses the connectivity of the entanglement wedge, $S_{R}$ quantifies the total mixed-state entanglement shared between the subsystems, including correlations inaccessible to entanglement negativity. In the backreacted AdS black hole background, this distinction becomes important: EWCS and consequently $S_{R}$ is enhanced by backreaction at high temperature whereas HEN exhibits a qualitatively different behavior in the low-temperature regime. This contrast highlights the complementary roles of EWCS, reflected entropy and negativity in characterizing the entanglement structure of strongly coupled plasmas.
@@ 4 Entanglement thermodynamics
$T=\frac{1}{4\pi z_{H}}\left(3-\frac{bz_{H}^{2}}{R^{2}}\right).$ (4.4)
@@
${\cal A}$ $=\frac{2LR^{2}}{z_{t}}\int_{0}^{1}\frac{1}{u^{2}\sqrt{1-u^{4}}}\left(1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}+\frac{bz_{H}^{2}}{R^{2}}\left(\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}-\frac{u^{2}z_{t}^{2}}{z_{H}^{2}}\right)\right)^{-\frac{1}{2}}du.$ (4.7)
We utilize the subsystem length and area formulas given in eq.˜4.6 and eq.˜4.7 to investigate the thermal and backreaction effects on HEE. For a field theory dual to a deformed AdS4 spacetime, the space of states is is characterized by both $T$ (via $z_{H}$) and $b$. This is in contrast to the Schwarzschild black hole in the AdS4 bulk scenario where $T$ is the unique scale controlling the high and low temperature limits of the entropy. In contrast, for the deformed black hole, the backreaction parameter $b$ also plays a role, as is evident from the requirement of temperature positivity in the bulk which follows from eq.˜4.4 as
@@ 4.1 Zero temperature
Thus, for deformed black holes at zero temperature, the small backreaction regime also corresponds to a large horizon radius $(l^{2}b\leq\frac{3R^{2}l^{2}}{z_{H}^{2}}<<1)$. With this condition, the lapse function from eq.˜4.2 simplifies
@@
${\cal A}$ $=\frac{2LR^{2}}{z_{t}}\int_{0}^{1}\frac{1}{u^{2}\sqrt{1-u^{4}}}\left(1-\frac{3u^{2}z_{t}^{2}}{z_{H}^{2}}+\frac{2u^{3}z_{t}^{3}}{z_{H}^{3}}\right)^{-\frac{1}{2}}du.$ (4.12)
@@
The first term in the expression for $\mathcal{A}$ corresponds to the divergent contribution which is same as of pure AdS4. We regularize the integral by imposing a UV cutoff $\epsilon=z_{b}/z_{t}$ and incorporate a counterterm $(-2LR^{2})/z_{b}$ to subtract the divergence. This procedure gives the following finite value for the area as
@@
$S_{A}^{finite}=\frac{1}{4G_{N}^{4}}\frac{LR^{2}}{l}\left[\frac{\pi\Gamma\left(\frac{3}{4}\right)\Gamma\left(-\frac{1}{4}\right)}{\Gamma\left(\frac{1}{4}\right)^{2}}+\frac{\pi\sqrt{2}\Gamma\left(\frac{5}{4}\right)b\,l^{2}}{3R^{2}\Gamma\left(\frac{3}{4}\right)^{3}}+\frac{16\pi\sqrt{2}l^{3}\Gamma\left(\frac{5}{4}\right)}{\Gamma\left(-\frac{1}{4}\right)^{3}z_{H}^{3}}+\mathcal{O}(\frac{l}{z_{H}})^{4}\right].$ (4.18)
Therefore, in the zero-temperature case, the HEE receives corrections that are linear and quadratic in $l$ and $l^{2}$. Notably, the emergence of a term linear in $l$ distinguishes this geometry from the extremal RN black hole case [17] where it is absent. By substituting the zero temperature mass $m_{0}=-R^{4}/z_{H}^{3}$ obtained from eq.˜4.3 and eq.˜4.9 into above eq.˜4.18, the renormalized HEE can be expressed in the following compact form
@@ 4.2 Low temperature regime
We now turn to the analysis of the subsystem $A$ in the boundary field theory dual to a deformed AdS4 black hole with small backreaction at low temperature. As follows from eq.˜4.8, when both the temperature and the backreaction are small, the horizon radius $z_{H}$ becomes large. In this limit, one has $bz_{H}^{2}/R^{2}\sim 1$ and $z_{H}$ sufficiently large such that $z_{t}\ll z_{H}$. Under these conditions, the lapse function $h(u)$ takes the form
@@ 4.3 First law of entanglement thermodynamics
$\Delta E_{A}=\int_{A}dx^{2}\ T_{tt}^{T\neq 0}-\int_{A}dx^{2}\ T_{tt}^{T=0}=\frac{Ll}{4\pi G_{N}^{4}R^{2}}(m-m_{0}),$ (4.32)
where $T_{tt}$ denotes the time component of the boundary stress-energy tensor and can be obtained using the holographic prescription of [4]. Now substituting the results from eq.˜4.31 and eq.˜4.32 into the first law like relation eq.˜4.30, the entanglement temperature is found to be
@@
We observe that the entanglement temperature $T_{\mathrm{ent}}$ is proportional to the inverse of the subsystem size. This universal scaling behavior is consistent with the entanglement thermodynamics of the asymptotically AdS geometries as discussed in [6]. This suggests that the small backreaction does not alter the fundamental scaling of the first law and the HEE continues to satisfy a well defined thermodynamic relation in the perturbative regime.
@@ 4.4 High temperature regime
In this section, we analyze the high temperature behavior of the HEE for a subsystem $A$ when the bulk deformed black hole has a small backreaction and is in the high temperature regime. From eq.˜4.8, it follows that in the limit of high temperature ($l/z_{H}>>1$) and small backreaction, the parameter $bz_{H}^{2}/3R^{2}$ becomes much smaller than unity i.e $\frac{bz_{H}^{2}}{3R^{2}}\ll 1$. By introducing the small parameter $\delta=bz_{H}^{2}/3R^{2}$, we can Taylor expand the lapse function $h(u)^{-1/2}$ around $\delta=0$ yielding
@@
$\mathcal{A}$ $=\frac{2LR^{2}}{z_{t}}\int_{0}^{1}\frac{1}{u^{2}\sqrt{1-u^{4}}}\left(\frac{1}{\sqrt{1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}}}+\frac{3\delta u^{2}z_{t}^{2}\left(1-\frac{uz_{t}}{z_{H}}\right)}{2z_{H}^{2}\left(1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}\right){}^{3/2}}\right)\,du.$ (4.36)
@@
As described in[31], for a static asymptotically AdS black hole, the extremal surface cannot penetrate the event horizon, necessitating that $z_{t}<z_{H}$. In the high temperature limit, however, the turning point approaches the horizon such that $z_{t}\to z_{H}$. By introducing a small positive parameter $\epsilon\ll 1$, we can parameterize this proximity as $z_{t}=(1-\epsilon)z_{H}$. Taking the limit $\epsilon\to 0$ in eq.˜4.39, the series converges to the following compact leading order expression as follows
@@
On substituting the form of $z_{H}$ from eq.˜4.4 in eq.˜4.40 leads to the following form of the parameter $\epsilon$ as
$\epsilon\approx\epsilon_{0}\,e^{-\frac{4\pi lT}{\sqrt{3}}(1+\delta)},$ (4.42)
where $\epsilon_{0}=\frac{e^{\sqrt{3}(C_{1}+\delta C_{2})}}{3}$.
In evaluating the area integral given in eq.˜4.36, divergent part is confined to the first term. Following [28], we isolate the divergence and express the remaining finite contribution as
${\cal A}^{finite}=\frac{2LR^{2}}{z_{t}}\int_{\frac{z_{b}}{z_{t}}}^{1}\frac{1}{u^{2}\sqrt{1-u^{4}}}\left(\frac{1}{\sqrt{1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}}}+\frac{3\delta u^{2}z_{t}^{2}\left(1-\frac{uz_{t}}{z_{H}}\right)}{2z_{H}^{2}\left(1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}\right){}^{3/2}}\right)\,du-\frac{2LR^{2}}{z_{b}}.$ (4.43)
@@
We can now use eq.˜4.38 in above eq.˜4.44 to simplify the extremal area as
@@
Here the term $s=\frac{R^{2}}{4G_{N}^{4}z_{H}^{2}}$ corresponds to the entropy density of the black hole as can be seen from from the eq.˜2.13. In eq.˜4.48, the first term follows a volume law scaling (proportional to the subsystem area in the $(2+1)$ dimensional dual field theory) and grows with temperature indicating that it represents the thermal part of the entropy. The subleading terms scale with the length of the entangling surface separating subsystem $A$ from its complement and thus represents entanglement contributions at high temperatures. Notably, the $\epsilon$-corrections exhibit an exponential decay with temperature analogous to the Schwarzschild-AdS case although they are modified by a small $\delta$ term arising from the backreaction.
@@
Entanglement negativity is a well established quantum information theoretic measure that captures the distillable quantum correlations in mixed states, making it particularly suitable for thermal or finite density systems where entanglement entropy alone fails to isolate genuine quantum correlations. Defined via the partial transpose of the reduced density matrix, negativity provides a computable upper bound on the distillable entanglement. Within the AdS/CFT framework, a holographic prescription for entanglement negativity in higher dimensions has been proposed, wherein suitable linear combinations of Ryu–Takayanagi surfaces encode the negativity of boundary mixed states [19, 37, 38, 54, 65, 1]. This construction allows one to extend the study of mixed state entanglement to strongly coupled holographic plasmas, and in what follows we analyze its behavior in the deformed AdS black hole background to compare with the trends observed for HEE.